# RiemannFormer: A Framework for Attention in Curved Spaces

### v1

@@ Abstract
This research endeavors to offer insights into unlocking the further potential of transformer-based architectures. One of the primary motivations is to offer a geometric interpretation for the attention mechanism in transformers. In our framework, the attention mainly involves metric tensors, tangent spaces, inner product, and how they relate to each other. These quantities and structures at discrete positions are intricately interconnected via the parallel transport of tangent vectors. To make the learning process more efficient, we reduce the number of parameters through ingenious predefined configurations. Moreover, we introduce an explicit mechanism to highlight a neighborhood by attenuating the remote values, given that transformers inherently neglect local inductive bias. Experimental results demonstrate that our modules deliver significant performance improvements relative to the baseline. More evaluation experiments on visual and large language models will be launched successively.
Keywords Transformer  $\cdot$ Self-Attention  $\cdot$ Curved Space  $\cdot$ Tangent Mapping  $\cdot$ Parallel Transport
@@ 1 Introduction
The transformer architecture, introduced in the seminal paper by Vaswani et al. [1], has revolutionized the field of natural language processing (NLP) and extended its impact to computer vision, audio processing, and other domains. The iterative application of transformer blocks allows tokens to progressively capture higher-level semantic concepts, making the architecture highly effective for tasks requiring contextual understanding.
@@
Since the self-attention mechanism is permutation-equivariant, it does not inherently capture the order of tokens. To address this, transformers incorporate positional encodings or positional embeddings, which are added to the input token embeddings before processing. Below, we provide a brief overview of positional encodings in transformer. Positional encodings are indispensable in transformer models, embedding sequential order into permutation-equivariant self-attention mechanisms to enable token position modeling in tasks like natural language processing and vision. The seminal sinusoidal positional encodings, introduced by Vaswani et al. [1], remain prevalent due to their deterministic nature and ability to generalize across arbitrary sequence lengths. In contrast, learned positional encodings, as used in BERT [2], optimize trainable embeddings for task-specific patterns but falter in extrapolating to longer sequences. Relative positional encodings, adopted in Transformer-XL [3], T5 [4], and TUPE [5], model token distance relationships, enhancing generalization for long-context tasks. Rotary Positional Embeddings (RoPE) efficiently encode relative positions by applying rotation matrices to query and key vectors [6], with extensions like xPos [7] incorporating dynamic scaling for stable long-sequence training. Recent innovations include ALiBi [8], which simplifies computation by introducing attention biases based on token distances. Prior methods like absolute and relative positional encoding are static post-training. Zheng et al. propose Data-Adaptive Positional Encoding (DAPE), which adjusts dynamically based on input context and learned priors [9]. In addition, Golovneva et al. introduced Contextual Position Encoding (CoPE) [10], a novel position encoding method that conditions positions on context by selectively incrementing positions based on model-determined tokens. Although methods like ALiBi and adaptive PEs reduce complexity, they may introduce instability in extremely long sequences, and their generalization across diverse domains remains underexplored.
@@
Since its adaptation to visual tasks with the Vision Transformer (ViT) [11], the transformer has demonstrated exceptional performance across a wide range of applications, from high-level tasks like classification [12] and detection [13] to low-level tasks such as super-resolution and generation. ViTs process images by dividing them into fixed-size patches, which are then treated as tokens similar to words in a sentence. These patches are embedded and fed into a transformer encoder, which leverages self-attention mechanisms to capture global relationships within the image. This approach allows ViTs to learn complex patterns and dependencies across the entire image, leading to impressive performance on various benchmarks. Positional encoding is a crucial component in the application of transformers to visual tasks, enabling the model to retain spatial information about the input image. Unlike natural language processing tasks where the sequence of words carries inherent positional information, images are divided into patches that lack such sequential context. Positional encodings allows the model to account for the order and arrangement of the patches, which is essential for understanding the spatial relationships and structure within the image. The original ViT model uses fixed positional encodings, similar to those used in the original Transformer model for NLP tasks. In addition to learnable and relative positional encodings, alternative schemes have emerged. RoPE excels in language models, particularly in enhancing transformer length extrapolation, yet its potential in computer vision remains largely unexplored. Heo et al. analyzes RoPE’s application to ViT, demonstrating significant improvements in visual tasks like classification, detection and segmentation, with minimal computational overhead [14]. Chu et al. introduce conditional positional encoding for ViT, dynamically generated based on local input token context [15]. These approaches potentially provide a more flexible and scalable approach to positional encoding in ViTs.
@@ 2 Methodology
Consider a sequence comprising $L$ input tokens, with their embeddings represented as $\{\boldsymbol{x}_{i}\}_{i=1}^{L}$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{D}$ denotes the $D$-dimensional embedding vector for the $i$-th token, devoid of positional information. In this paper, we adopt an alternative perspective: we posit the existence of a curved space $\mathcal{M}$ wherein the encoding of a token sequence corresponds to identifying a sequence of points $\{\mathbf{p}_{i}\}$, or a grid of points $\{\mathbf{p}_{ij}\}$ for image data.
@@
The geometric structure of a curved space implicitly supports positional information for sequences or grids of points. However, in such a non-flat space, vectors in tangent spaces at distinct points cannot be directly operated upon. What strategies can be employed to effectively tackle this challenge? Specifically, the vectors $\mathbf{q}_{i}$ and $\mathbf{k}_{i}$ reside in the tangent spaces $T_{\mathbf{p}_{i}}\mathcal{M}$ at position $\mathbf{p}_{i}$. For instance, $\mathbf{q}_{m}$ and $\mathbf{k}_{m}$ both belong to the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$, whereas $\mathbf{q}_{m}$ and $\mathbf{k}_{n}$ are situated in the tangent spaces at points $\mathbf{p}_{m}$ and $\mathbf{p}_{n}$, respectively. The inner product in the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$ is defined as $\mathbf{x}_{1}^{\top}\mathbf{M}_{m}\mathbf{x}_{2}$, where $\mathbf{M}_{m}$ is a Riemannian metric specifying the inner product structure at $\mathbf{p}_{m}$. Since $\mathbf{k}_{n}$ does not belong to the tangent space at $\mathbf{p}_{m}$, it is not appropriate to directly perform inner products such as $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathbf{k}_{n}$. To address this, a transformation $\mathtt{f}_{n,m}(\mathbf{k}_{n})$ is required to appropriately transport $\mathbf{q}_{n}$ from the tangent space at $\mathbf{p}_{n}$ to that at $\mathbf{p}_{m}$. Only after this transformation can a valid inner product be executed, given by $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathtt{f}_{n,m}(\mathbf{k}_{n})$.
@@
2) Consider a Riemannian manifold $\mathcal{M}$ equipped with its associated Riemannian connection $\nabla$. Let $\mathbf{V}$ be a parallel vector field along a parameterized curve $c(t):[0,1]\to\mathcal{M}$. The vector field $\mathbf{V}$ maintains a constant magnitude along the curve $c$.
@@
Leveraging the second property mentioned above, we derive the following equation:
$\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}=\mathcal{P}_{n\rightarrow m}({\bf k}_{n})^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}({\bf k}_{n})$ (1)
First, we introduce a common reference space $\mathbb{R}^{D}$, upon which the linear transformations between tangent spaces at various points are subsequently learned. The linear transformation to be learned at point $\mathbf{p}_{i}$ is denoted by $\mathcal{P}_{i}:\mathbb{R}^{D}\to T_{\mathbf{p}_{i}}\mathcal{M}$ and can be represented by a $D\times D$ matrix $\mathbf{T}_{i}$. In this study, we term the transformation $\mathcal{P}_{i}^{-1}$ as tangent mapping or tangent aligning.
Building upon the foregoing analysis, we have
@@
By substituting this result into Eq. (1), the following equation can be derived:
$\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}=\mathbf{k}_{n}^{\top}\left(\mathbf{T}^{-1}_{n}\right)^{\top}\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}$ (3)
@@
The above relationship of these matrices is what we must consider when predefining them.
To clarify again, the aim of this work is to enable the model to learn the geometric structure of the curved space where the sequence data reside. The primary determinants are the metric tensors and the transformations between the tangent spaces at the positions involved. Given the intricate nature and the extensive array of parameters involved, it is essential to establish predefined configurations for these matrices.
A Riemannian metric $\mathbf{M}_{i}$ can be represented by a symmetric positive definite matrix, but for the sake of simplifying the problem, we preset it as a diagonal matrix:
$\mathbf{M}_{m}=\begin{bmatrix}s_{1}^{m}&&&\\ &s_{2}^{m}&&\\ &&\ddots&\\ &&&s_{D}^{m}\end{bmatrix},$ (6)
where $s_{i}>0$ are learnable parameters. To further reduce the number of parameters, a simpler presetting, a scalar matrix $\mathbf{M}_{m}=s^{m}\mathbf{I}$ can be employed, where $\mathbf{I}$ is the identity matrix. This is also known as a conformal deformation of $\mathbb{R}^{D}$. To impose the constraint $s<1$, we may proceed by defining it as $s=\frac{\exp(w)}{\exp(w)+\alpha}$, where $w$ is a learnable parameter and $\alpha$ denotes a hyperparameter. Alternatively, one might opt for $s=\exp(w)$, which permits $s>1$. For long sequences, we can decelerate the rate of exponential growth, by setting the parameter $s^{\log_{\beta}(m)}$. In addition, the learnable metric matrix of general form is given in the Appendix section.
@@
To facilitate the model to effectively learn these linear transformations, we first decompose them into two components, including scaling and length-preserving in $\mathbb{R}^{D}$. The introduction of scaling part is motivated by the characteristics of the curved space and the existence of Riemannian metric. The scaling part is represented by a diagonal matrix or a scalar matrix $s\mathbf{I}$. An orthogonal matrix is used for length-preserving, and a rotation is naturally a candidate. A rotation in $n$-dimensional Euclidean space is a linear transformation that preserves lengths of vectors, and it can be represented by an $D\times D$ orthogonal matrix in the special orthogonal group $SO(D)$. It is well known that an element of $SO(D)$ can be generated from the exponential of a skew-symmetric matrix $\mathbf{X}$, denoted as $\exp(\mathbf{X})$.

### v2

@@ Abstract
This research endeavors to unlock the further potential of transformer-based architectures by providing a geometric interpretation of the self-attention mechanism. In our framework, we reframe attention as interactions on a potential Attention Manifold. Queries and keys are treated as vectors within the manifold’s intrinsic geometry, residing in the tangent bundle, where the geometry of each tangent space is governed by a Riemannian metric. This metric is the fundamental object that defines both the local inner product for similarity computation and the connection that determines parallel transport. In stark contrast, we propose that value vectors are sections of a separate, trivial vector bundle, decoupling the geometric alignment of $Q/K$ from the feature aggregation of $V$. Our primary contribution is a novel, first-principles construction of this parallel transport using Lie group theory, where for 2D data, we derive a positional encoding from the exceptional isomorphism $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$. This allows us to factorize 4D rotations into two commuting isoclinic rotations aligned with the spatial dimensions. This dual-bundle structure provides a principled foundation for our explicit mechanism that highlights a neighborhood by attenuating remote values. Experimental results demonstrate that our modules deliver significant performance improvements over strong baselines, and further evaluations on visual and large language models are underway.
Keywords Transformer $\cdot$ Self-Attention $\cdot$ Curved Space $\cdot$ Tangent Bundle $\cdot$ Fiber Bundle $\cdot$ Parallel Transport
@@ 1 Introduction
The transformer architecture, introduced in the seminal paper by Vaswani et al. [Vaswani2017], has revolutionized the field of natural language processing (NLP) and extended its impact to computer vision, audio processing, and other domains. The iterative application of transformer blocks allows tokens to progressively capture higher-level semantic concepts, making the architecture highly effective for tasks requiring contextual understanding.
@@
Since the self-attention mechanism is permutation-equivariant, it does not inherently capture the order of tokens. To address this, transformers incorporate positional encodings or positional embeddings, which are added to the input token embeddings before processing. Below, we provide a brief overview of positional encodings in transformer. Positional encodings are indispensable in transformer models, embedding sequential order into permutation-equivariant self-attention mechanisms to enable token position modeling in tasks like natural language processing and vision. The seminal sinusoidal positional encodings, introduced by Vaswani et al. [Vaswani2017], remain prevalent due to their deterministic nature and ability to generalize across arbitrary sequence lengths. In contrast, learned positional encodings, as used in BERT [DevlinCLT19], optimize trainable embeddings for task-specific patterns but falter in extrapolating to longer sequences. Relative positional encodings, adopted in Transformer-XL [Dai2019TransformerXLAL], T5 [2020t5], and TUPE [KeHL21], model token distance relationships, enhancing generalization for long-context tasks. Rotary Positional Embeddings (RoPE) efficiently encode relative positions by applying rotation matrices to query and key vectors [Su2024RoPE], with extensions like xPos [sun2023] incorporating dynamic scaling for stable long-sequence training. Recent innovations include ALiBi [alibi], which simplifies computation by introducing attention biases based on token distances. Prior methods like absolute and relative positional encoding are static post-training. Zheng et al. propose Data-Adaptive Positional Encoding (DAPE), which adjusts dynamically based on input context and learned priors [zheng2024]. In addition, Golovneva et al. introduced Contextual Position Encoding (CoPE) [golovneva2024], a novel position encoding method that conditions positions on context by selectively incrementing positions based on model-determined tokens. Although methods like ALiBi and adaptive PEs reduce complexity, they may introduce instability in extremely long sequences, and their generalization across diverse domains remains underexplored.
@@
Since its adaptation to visual tasks with the Vision Transformer (ViT) [DosovitskiyB0WZ21], the transformer has demonstrated exceptional performance across a wide range of applications, from high-level tasks like classification [DeiT2021] and detection [Carion2020] to low-level tasks such as super-resolution and generation. ViTs process images by dividing them into fixed-size patches, which are then treated as tokens similar to words in a sentence. These patches are embedded and fed into a transformer encoder, which leverages self-attention mechanisms to capture global relationships within the image. This approach allows ViTs to learn complex patterns and dependencies across the entire image, leading to impressive performance on various benchmarks. Positional encoding is a crucial component in the application of transformers to visual tasks, enabling the model to retain spatial information about the input image. Unlike natural language processing tasks where the sequence of words carries inherent positional information, images are divided into patches that lack such sequential context. Positional encodings allows the model to account for the order and arrangement of the patches, which is essential for understanding the spatial relationships and structure within the image. The original ViT model uses fixed positional encodings, similar to those used in the original Transformer model for NLP tasks. In addition to learnable and relative positional encodings, alternative schemes have emerged. RoPE excels in language models, particularly in enhancing transformer length extrapolation, yet its potential in computer vision remains largely unexplored. Heo et al. analyzes RoPE’s application to ViT, demonstrating significant improvements in visual tasks like classification, detection and segmentation, with minimal computational overhead [heo2024ropevit]. Chu et al. introduce conditional positional encoding for ViT, dynamically generated based on local input token context [chu2023CPVT]. These approaches potentially provide a more flexible and scalable approach to positional encoding in ViTs.
@@ 2 Methodology
Consider a sequence comprising $L$ input tokens, with their embeddings represented as $\{\mathbf{x}_{i}\}_{i=1}^{L}$, where $\mathbf{x}_{i}\in\mathbb{R}^{D}$ denotes the $D$-dimensional embedding vector for the $i$-th token, devoid of positional information. In this paper, we adopt an alternative perspective: we posit the existence of a curved space $\mathcal{M}$ wherein the encoding of a token sequence corresponds to identifying a sequence of points $\{\mathbf{p}_{i}\}$, or a grid of points $\{\mathbf{p}_{ij}\}$ for image data, etc.
2.1 Attention Manifold
We introduce the concept of the Attention Manifold, a novel framework for interpreting Transformer self-attention through non-Euclidean geometry. In this view, token positions are points on a manifold, and their feature embeddings are vectors in the corresponding tangent spaces. The manifold’s geometry is defined by a Riemannian metric, which modulates local interactions and encodes semantic saliency, and a parallel transport operator, which ensures consistent comparisons between vectors at different positions. This geometric formulation provides a first-principles approach to designing attention mechanisms with strong, built-in inductive biases.
Our approach is motivated by a fundamental shift in perspective from an extrinsic to an intrinsic geometric viewpoint. From this standpoint, a manifold is not defined by the absolute coordinates of its points within some ambient space, but rather by the complete set of relationships and transformations that connect them. Consequently, our primary objective is not to parameterize the absolute locations of tokens. Instead, we focus entirely on the crucial task: to directly define a consistent and computationally tractable protocol for the transport of tangent vectors (the feature embeddings) between the tangent spaces at discrete points. In essence, what matters is not where the tokens are, but how their representations can be transformed into one another.
Formally, in differential geometry, the parallel transport of a vector field V{V} along a parameterized curve $c(t)$ is governed by the condition that its covariant derivative along the curve is zero. In local coordinates, this translates to a system of differential equations:
$\frac{dv^{k}}{dt}+\Gamma^{k}_{ij}\frac{dc^{i}}{dt}v^{j}=0,$
where $c(t)$ is the path between points, and $\Gamma^{k}_{ij}$ are the connection coefficients (e.g., Christoffel symbols) that encode the manifold’s curvature.
@@ 2.1 Attention Manifold
However, this formal definition presents a fundamental challenge for our discrete and non-parametric setting. We intentionally avoid the complexity and high cost of explicitly defining the manifold’s embedding, the continuous paths $c(t)$ between tokens, or the connection coefficients $\Gamma^{k}_{ij}$. Consequently, this system of equations, while theoretically fundamental, is practically intractable within our framework as we lack the necessary components for its solution. This raises the central question: how can we define a consistent and computationally tractable parallel transport operator without recourse to the underlying continuous geometry? Specifically, the vectors $\mathbf{q}_{i}$ and $\mathbf{k}_{i}$ reside in the tangent spaces $T_{\mathbf{p}_{i}}\mathcal{M}$ at position $\mathbf{p}_{i}$. For instance, $\mathbf{q}_{m}$ and $\mathbf{k}_{m}$ both belong to the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$, whereas $\mathbf{q}_{m}$ and $\mathbf{k}_{n}$ are situated in the tangent spaces at points $\mathbf{p}_{m}$ and $\mathbf{p}_{n}$, respectively. The inner product in the tangent space $T_{\mathbf{p}_{m}}\mathcal{M}$ at point $\mathbf{p}_{m}$ is defined as $\mathbf{x}_{1}^{\top}\mathbf{M}_{m}\mathbf{x}_{2}$, where $\mathbf{M}_{m}$ is a Riemannian metric specifying the inner product structure at $\mathbf{p}_{m}$. Since $\mathbf{k}_{n}$ does not belong to the tangent space at $\mathbf{p}_{m}$, it is not appropriate to directly perform inner products such as $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathbf{k}_{n}$. To address this, a transformation $\mathtt{f}_{n,m}(\mathbf{k}_{n})$ is required to appropriately transport $\mathbf{q}_{n}$ from the tangent space at $\mathbf{p}_{n}$ to that at $\mathbf{p}_{m}$. Only after this transformation can a valid inner product be executed, given by $\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathtt{f}_{n,m}(\mathbf{k}_{n})$.
Let us now formalize the attention mechanism within our geometric framework. We consider a discrete sequence of points, $\{\mathbf{p}_{i}\}_{i=1}^{N}$, defined on the Attention Manifold $\mathcal{M}$. The collection of all tangent spaces on the manifold, $\{T_{\mathbf{p}}\mathcal{M}\}_{\mathbf{p}\in\mathcal{M}}$, forms a structure known as the Tangent Bundle, denoted $T\mathcal{M}$. We conceptualize the query and key representations as sections of this tangent bundle. The query and key vector fields, $Q$ and $K$, therefore assign a tangent vector $\mathbf{q}_{i},\mathbf{k}_{i}\in T_{\mathbf{p}_{i}}\mathcal{M}$ to each point $\mathbf{p}_{i}$. These vectors live within the intrinsic, and potentially curved, geometry of the manifold. In stark contrast, the value representations are treated as sections of a Trivial Vector Bundle, $\mathcal{M}\times\mathbb{R}^{D_{v}}$. This construction places the value vectors $\{\mathbf{v}_{i}\}$ in a global, geometry-agnostic Euclidean space. This trivial bundle construction formally decouples the geometric interactions of queries and keys on the manifold from the value representations. A principal advantage of this decoupling is the architectural flexibility it affords: the dimensionality of the value vectors, $D_{v}$, is not required to match the dimensionality of the tangent spaces, $D$, where queries and keys reside. This allows for independent design choices, mirroring the flexibility found in standard Multi-Head Attention architectures. At each point $\mathbf{p}_{i}\in\mathcal{M}$, we sample a specific query and key vector $\mathbf{q}_{i}\in T_{\mathbf{p}_{i}}\mathcal{M}$ and $\mathbf{k}_{i}\in T_{\mathbf{p}_{i}}\mathcal{M}$ from their respective fields, where the tangent space has dimension $D$. The core of the attention mechanism is the interaction between a query vector $\mathbf{q}_{i}$ at point $\mathbf{p}_{i}$ and key vectors from all other points, $\{\mathbf{k}_{j}\}$. To perform a geometrically consistent comparison, each key vector $\mathbf{k}_{j}$ is first transported from its native tangent space $T_{\mathbf{p}_{j}}\mathcal{M}$ to the query’s tangent space $T_{\mathbf{p}_{i}}\mathcal{M}$. The attention scores are then computed via the local inner product.
Finally, these scalar scores are normalized to form attention weights. These weights are then used to compute a weighted sum of the value vectors $\{\mathbf{v}_{j}\}$ in their common Euclidean space $\mathbb{R}^{D_{v}}$. This aggregation, which does not require parallel transport of the values, yields the updated representation at position $\mathbf{p}_{i}$, which itself is a vector in $\mathbb{R}^{D_{v}}$.
2.2 Parallel Transport
@@
2) Consider a Riemannian manifold $\mathcal{M}$ equipped with its associated Riemannian connection $\nabla$. Let V{V} be a parallel vector field along a parameterized curve $c(t):[0,1]\to\mathcal{M}$. The vector field V{V} maintains a constant magnitude along the curve $c$.
@@
To formalize the geometric concepts of vector magnitude and the defining property of our parallel transport operator, we can express the squared norm of a tangent vector in both abstract and computational forms. A tangent vector, such as the key vector $\mathbf{k}_{n}$ at position $\mathbf{p}_{n}$, is expressed in a local coordinate basis $\{\partial_{i}\}$ as a linear combination $\sum_{i}k^{i}_{n}\partial_{i}$, where $\{k^{i}_{n}\}$ are its components.
The squared norm of this vector is computed via the inner product defined by the metric tensor $\mathbf{M}_{n}$ at that point. This relationship is articulated by the following equation:
$\underbrace{\langle\sum\nolimits_{i}k^{i}_{n}\partial_{i},\sum\nolimits_{j}k^{j}_{n}\partial_{j}\rangle_{\mathbf{p}_{n}}}_{\text{Abstract Inner Product}}=\underbrace{\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}}_{\text{Computational Form}}=\underbrace{\mathcal{P}_{n\rightarrow m}({\bf k}_{n})^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}({\bf k}_{n})}_{\text{Isometry under Parallel Transport}}.$ (1)
By the property of bilinearity, the inner product expands to $\sum_{i,j}k^{i}_{n}k^{j}_{n}\langle\partial_{i},\partial_{j}\rangle_{\mathbf{p}_{n}}$. The components of the metric tensor $\mathbf{M}_{n}$ are, by definition, the inner products of the basis vectors themselves: $(\mathbf{M}_{n})_{ij}=\langle\partial_{i},\partial_{j}\rangle_{\mathbf{p}_{n}}$. Substituting this definition directly yields the familiar quadratic form $\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}$, where $\mathbf{k}_{n}$ is the column vector of components $\{k^{i}_{n}\}$. This term represents the squared magnitude of the vector $\mathbf{k}_{n}$ according to the local geometry at $\mathbf{p}_{n}$.
(a) Maps between tangent spaces.
First, we introduce a common reference space $\mathbb{R}^{D}$, upon which the linear transformations between tangent spaces at various points are subsequently learned. As shown in Fig. (1(a)), the linear transformation to be learned at point $\mathbf{p}_{i}$ is denoted by $\mathcal{P}_{i}:\mathbb{R}^{D}\to T_{\mathbf{p}_{i}}\mathcal{M}$ and can be represented by a $D\times D$ matrix $\mathbf{T}_{i}$. In this study, we term the transformation $\mathcal{P}_{i}^{-1}$ as tangent mapping or tangent aligning.
Building upon the foregoing analysis, we formally define the parallel transport operator, $\mathcal{P}_{n\rightarrow m}$, which transports a vector $\mathbf{k}_{n}$ from the tangent space at position $n$ to that at position $m$. This is achieved by composing the inverse transformation at the source with the forward transformation at the destination:
@@
This formulation inherently satisfies the property of transitivity. A composition of transport operations from $n$ to $m$ and subsequently from $m$ to $l$ simplifies to a single direct transport from $n$ to $l$. This is demonstrated by the algebraic cancellation of the intermediate operators, which confirms the robustness of our framework and a consistent geometric structure:
$\mathcal{P}_{n\rightarrow m\rightarrow l}({\bf k}_{n})$ $=\mathcal{P}_{l}\circ\mathcal{P}^{-1}_{m}\left(\mathcal{P}_{m}\circ\mathcal{P}^{-1}_{n}({\bf k}_{n})\right)$ (3)
$=\mathbf{T}_{l}\mathbf{T}^{-1}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}$
$=\mathbf{T}_{l}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}$
$=\mathcal{P}_{n\rightarrow l}({\bf k}_{n}).$
Substituting Eq. (2) into Eq. (1), we obtain:
$\mathbf{k}_{n}^{\top}\mathbf{M}_{n}\mathbf{k}_{n}=\mathbf{k}_{n}^{\top}\left(\mathbf{T}^{-1}_{n}\right)^{\top}\mathbf{T}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}^{-1}_{n}\mathbf{k}_{n}.$ (4)
@@
The above relationship of these matrices is what we must consider when predefining them. The Riemannian metric $\mathbf{M}_{i}$ is a symmetric positive matrix. To facilitate the model to effectively learn the linear transformations $\mathbf{T}_{i}$, we first decompose them into two components, including scaling and length-preserving in $\mathbb{R}^{D}$: $\mathbf{T}_{i}=\mathbf{S}_{i}\mathbf{R}_{i}$. The introduction of scaling part $\mathbf{S}_{i}$ is motivated by the characteristics of the curved space and the existence of Riemannian metric. The scaling part is represented by a symmetric positive matrix, and its simple form is a diagonal matrix or a scalar matrix $s\mathbf{I}$, where $\mathbf{I}$ is the identity matrix. For instance, we set $\mathbf{S}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}$, then we have $\mathbf{T}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}\mathbf{R}_{i}$. $\mathbf{R}_{i}$ is an orthogonal matrix used for length-preserving, and a rotation is naturally a candidate. A rotation in $n$-dimensional Euclidean space is a linear transformation that preserves lengths of vectors, and it can be represented by an $D\times D$ orthogonal matrix in the special orthogonal group $\mathbf{SO}(D)$.
Given these definitions, it is straightforward to verify that Eq. (6) holds. Furthermore, the metric-induced inner product can be readily converted into the standard inner product as shown in the following formula, thereby simplifying its implementation.
$=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{M}_{m}^{-\frac{1}{2}}\mathbf{R}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$
$=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}^{\frac{1}{2}}\mathbf{R}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}$
To reiterate, the central aim of our work is to empower the model to learn the intrinsic geometry of the Attention Manifold where sequence data reside. This geometry is fully characterized by the fundamental objects: the set of Riemannian metric tensors $\{\mathbf{M}_{m}\}$ that define local inner products and the transformations between tangent spaces. However, a fully general and learnable parameterization for these objects would be computationally intractable and parametrically prohibitive. Therefore, the core of our contribution lies in introducing a principled, parameter-efficient construction for these geometric structures, using predefined configurations derived from Lie group theory.
A Riemannian metric $\mathbf{M}_{m}$ can be represented by a symmetric positive definite matrix, but for the sake of simplifying the problem, we preset it as a symmetric positive definite diagonal matrix:
$\mathbf{M}_{m}=\operatorname{diag}{(s^{(1)}_{m},s^{(2)}_{m},\cdots,s^{(D)}_{m})}$, (8)
where $\langle\partial_{i},\partial_{i}\rangle_{\mathbf{p}_{m}}=s^{(i)}_{m}$, the superscript $(i)$ denotes the component index. To enhance computational efficiency and reduce parametric complexity, we may constrain the metric tensor $\mathbf{M}_{m}$ to be a scalar matrix of the form $\mathbf{M}_{m}=s_{m}\mathbf{I}$, where $s_{m}>0$. This choice is geometrically equivalent to inducing a conformal deformation on the standard Euclidean space $\mathbb{R}^{D}$. In our geometric framework, this matrix $\mathbf{M}_{m}$ is not merely a scaling factor; it is interpreted as the metric tensor that defines the local geometry of the tangent space at position $m$. Its primary function is to modulate the inner product, thereby assigning a notion of semantic importance or saliency to the token at that position. The standard self-attention mechanism implicitly assumes a uniform Euclidean metric ($\mathbf{M}=\mathbf{I}$), treating all positions as equally important a priori. By introducing a position-dependent metric $\mathbf{M}_{m}$, we allow the model to differentiate the intrinsic significance of tokens. A larger scalar $s_{m}$ effectively amplifies the length of the query vector $\mathbf{q}_{m}$ and, consequently, the magnitude of its computed attention scores to all other keys. Thus, $s_{m}$ acts as a control knob for the token’s influence: a higher value signifies greater semantic importance, causing its query to have a stronger impact on the overall attention distribution.
While the general formulation $\mathbf{M}_{m}=s_{m}\mathbf{I}$ allows for arbitrary, position-specific importance, it does not explicitly encode a structural bias related to sequential order. To introduce such a prior, we can parameterize the scaling factor as an exponential function of the position index $m$. This is elegantly achieved by setting $\mathbf{M}_{m}=s^{m}\mathbf{I}$, where $s>0$ is a single, globally shared, learnable parameter. This highly parsimonious choice allows the model to learn a global decay rate ($s<1$) or a growth rate ($s>1$) of semantic importance, thereby systematically weighting tokens based on their absolute position in the sequence. We refer to this scheme as Conformal Exponential Scaling (CES). To impose the constraint $s<1$, we may proceed by defining it as $s=\frac{\exp(w)}{\exp(w)+\alpha}$, where $w$ is a learnable parameter and $\alpha$ denotes a hyperparameter. Alternatively, one might opt for $s=\exp(w)$, which permits $s>1$. For long sequences, we can decelerate the rate of exponential growth, by setting the parameter $s^{\log_{\beta}(m)}$. In addition, the learnable metric matrix of general form is given in the Appendix section.
2.3 A Principled Decomposition of High-Dimensional Rotations
A rotation in $\mathbb{R}^{D}$ can be represented by a $D\times D$ matrix in the special orthogonal group $\mathbf{SO}(D)$. It is well known that an element of $\mathbf{SO}(D)$ can be generated from the exponential of a skew-symmetric matrix $X$, denoted as $\exp(X)$. Specifically, the orthogonal Lie algebra $\mathfrak{so}(D=2k)$:
$\{X\mid X^{\top}+X=0\}$
yields $D$-dimensional special orthogonal matrices via the exponential map: $\exp(X)\in\mathbf{SO}(D)$.